OpenStudy (anonymous):

There are two complementary angles. The larger angle is 15 degrees more than twice the smaller angle. Find the measure of each angle.

OpenStudy (anonymous):

Let the smaller angle be = x degrees Then larger angle is 2x + 15 degrees

OpenStudy (anonymous):

Since they r complementary angles, theri sum must be 90 degrees so x + (2x + 15) = 90

OpenStudy (anonymous):

Therefore 3x + 15 = 90 3x = 90 - 15 x = 75/3 x = 25

OpenStudy (anonymous):

Hence smaller angle is 25 degrees The larger angle is 2 x 25 + 15 =65 degrees Ok??

OpenStudy (anonymous):

ok, thank you so much

OpenStudy (anonymous):

You are welcome. Got a medal for me??

OpenStudy (anonymous):

how do i do the medal thing?

OpenStudy (anonymous):

Just go to the first post from me and click on the button "Give Medal"

OpenStudy (anonymous):

:), ok gave you a medal